[Hallc_running] [New Logentry] Hall C RC daily update
simona at jlab.org
simona at jlab.org
Tue Aug 4 15:55:01 EDT 2020
Last night Dave and Bill finished up the Moller measurement around 8 pm: https://logbooks.jlab.org/entry/3820234
After the Moller measurement the reference cell Ion Chamber calibration was performed which allowed us to take our first production run on a glass cell. Due to beam trips we were limited to 15 muA (out of the desired 30 muA): https://logbooks.jlab.org/entry/3820411
We didn't get much further than that due to TS Isaias: https://logbooks.jlab.org/entry/3820368 , https://logbooks.jlab.org/entry/3820432 , https://logbooks.jlab.org/entry/3820536
The machine is still down and MCC is working to bring everything back up.
After noon Hall C has been surveyed and then put in Restricted access. We are ready for the sweep whenever MCC has time to support it (they will be very busy for a while working to bring the machine back up).
The target experts have been working during the day shift to improve the target polarization.
Once beam is restored to Hall C we will allow up to 45 muA rastered CW beam to the dump (with the target ladder in the "no target" position and centered beam) to verify that stable 30 muA can be delivered. If that's successful, we will do a BCM calibration (needs up to 42 muA) and then move to production on 3He. If we are limited to less than 30 muA, we will take production data and prorate the ABUs.
We will have an RC meeting today at 4:15 pm. This is how to connect via bluejeans: https://bluejeans.com/861439920
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3820548
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running